OpenStudy (anonymous):

Ten is the geometric mean of 4 and what number? 2 40 25 8

OpenStudy (anonymous):

Gm^2=a*b c^2=4*b 10^2=4*b b=10^2/4=25

OpenStudy (anonymous):

What is the geometric mean of 3 and 15? 45 22.5

OpenStudy (anonymous):

U try it no

OpenStudy (anonymous):

idk how to do it

OpenStudy (anonymous):

th geometric mean of a and b is \(\large \sqrt{ab} \)